OpenStudy (anonymous):

how can i solve this system of equation 3x+8=2 7x+3y=1

OpenStudy (radar):

Solve for x in the first equation 3x+8=2 after subtracting 8 and dividing both sides by 3, getting x=-6/3=-2. Now substitute -2 for x in the second equation. Then solve for y. Is that enough help?

OpenStudy (anonymous):

AAAWW THNX IJUS NEEDD TO SEE HOW ITS SOLVED ND THEN MAYBE ICOULD TAKE IT FROM THERE THNX HELPED:)

OpenStudy (anonymous):

HEY IGOTTA QUESTION THO

OpenStudy (radar):

Fine, when you substitute the -2 for x in the second equation you get: 7(-2)+3y=1 or -14+3y=1 Now solve for y.

OpenStudy (radar):

Go ahead with the question.

OpenStudy (anonymous):

SO IM GOING TO HAVE TWO ANSWERS FOR THIS EQUATION WHICH IS X AND Y

OpenStudy (radar):

X was solved in my first response. X is -2 Now we will solve for Y. Using the second equation substitute (-2) for the x value. This will get you 7(-2)+3y=1 or -14+3y=1 Add 14 to both sides getting 3y=15. Now solve for y by dividing both sides by 3 getting y=5. Now you have the value of both x and y.

OpenStudy (anonymous):

OK COOL THATS THE ANSWER THAT I GOT THANK YU

OpenStudy (radar):

O.K. Hang in there. It will make more sense with practice.

OpenStudy (anonymous):

OKAY